OpenStudy (anonymous):

inductive reasoning pattern 1,2,4,7,11,16,22,29 5,15,45,135,405,1,215 10,20,11,18,12,16,13,14,14,12,15 1,4,9,16,24,31,34,39,46 i need help with finding pattern with these

OpenStudy (anonymous):

Find the difference between each set of 2 numbers. What's the difference between 1 and 2? Between 2 and 4? Between 4 and 7?

OpenStudy (anonymous):

oh i get it...

OpenStudy (anonymous):

+1 to each number.. Example: 1+1=2, 2+2=4, 4+3=7, 7+4=11 .. catching my drift ?

OpenStudy (anonymous):

it was easy..

OpenStudy (anonymous):

yup exactly @mlovegurah Good!

OpenStudy (anonymous):

lol thanks @PigPresident

OpenStudy (anonymous):

yup hehe